# Multi-Robot Navigation in Social Mini-Games: Definitions, Taxonomy, and Algorithms

### v3

@@ 4.1.2 Right-Hand-Rule (RHS) [64, 147]:
RHS implements perturbation following the right hand rule (clockwise priority) to resolve conflicts among agents in SMGs. The idea behind RHS is to enforce an ordering to agent’s movements, avoiding strictly random paths that can lead to deadlocks, especially in crowded or tight spaces. RHS is designed to reduce the chances of deadlocks caused by overly structured and predictable paths. This rule is controlled by barrier functions, ensuring that even with noisy movements, agents still follow safety and operational rules. The nature of this method allows agents to respond dynamically to changing scenarios in real time. However, due to the rule enforcement, this method brings some limitations as well. While, on the one hand, it aids in resolving the deadlocks and creating smooth navigation in shared spaces, on the other hand, it can violate fairness constraints resulting in less efficient routes, highlighting a trade-off between goal completion and optimality.
4.1.3 Auction-Based [85, 131]:
@@
The Auction-Based algorithm combines Control Barrier Functions (CBF) in a auction based manner for prioritizing the agents for motion planning. By creating a competitive environment among agents and carefully adjusting their speeds, the Auction-Based method aims to improve navigation efficiency, even in crowded scenarios. A key feature of this method is its ability to prioritize agents. Through velocity scaling, it creates a hierarchy where some agents are given priority, leading to more efficient navigation preventing deadlocks, especially in congested areas. However, while velocity scaling helps in managing priority-based navigation, it can sometimes increase the time it takes for certain agents to reach their destinations. This can potentially lengthen the overall navigation time, reflecting a trade-off in the prioritization scheme used by the Auction-Based method. Despite this trade-off, the method aims for global optimization in path planning, ensuring more orderly and efficient navigation, especially in scenarios with many agents.
4.1.4 IMPC-DR [77]:
@@
IMPC-DR is a method specifically designed to avoid deadlocks using infinite-horizon Model Predictive Control (MPC). The MPC uses a special structure called modified buffered Voronoi [159] with a warning band. This structure helps to understand when deadlocks might happen, similar to a state of force equilibrium. When a potential deadlock is detected, the method quickly uses an adaptive resolution scheme to resolve it. This ensures that no stable deadlocks can occur under specific conditions, allowing smooth navigation for the multi-robot system. The planning algorithm in this method ensures that the optimization problem remains feasible at every time step while meeting input and model constraints. It works simultaneously across all robots, promoting coordinated and synchronized navigation. The method relies on local communication, avoiding the need for a central hub or extensive networks, which ultimately reduces the communication burden and improves the scalability and real-time response. Local communication also strengthens the decentralized nature of the method, making it a reliable solution to generate collision-free trajectories in shared spaces. This communication approach systematically prevents and resolves deadlocks in multi-robot navigation.
4.1.5 ORCA-MAPF [134, 46]:
@@
ORCA-MAPF (Optimal Reciprocal Collision Avoidance - Multi-Agent Path Finding) addresses the challenge of preventing collisions in multi-agent navigation, especially in decentralized settings with limited communication and sensing. In such environments, collision avoidance is usually reactive, based on local observations or sparse communications among agents. While strategies like ORCA are known for their efficiency and scalability, they often struggle in complex scenarios, such as navigating through narrow passages. In these cases, deadlocks are common due to the self-centered behavior of agents, preventing them from reaching their goals. The innovative approach of ORCA-MAPF uses locally confined multi-agent path finding (MAPF) solvers to coordinate sub-groups of agents at risk of deadlock. This method includes a simple and creates a grid-based MAPF instance for modern MAPF solvers. The grid-based representation provides a structured environment, making the problem easier to solve. ORCA-MAPF significantly improves the success rate of navigation in simple or empty shared spaces, increasing the safety rate from 15% to 99% in some tests. This improvement highlights ORCA-MAPF’s effectiveness in handling deadlocks and enhancing overall navigation success in decentralized multi-agent systems. By integrating MAPF solvers, ORCA-MAPF overcomes the limitations of existing collision avoidance techniques, offering more robust and reliable multi-agent navigation in complex, real-world environments. Through strategic coordination and effective deadlock management, ORCA-MAPF advances the field of multi-agent navigation, especially in decentralized and communication-limited settings.
4.2 Comparative Analysis in SMGLib
SMGLib is a simulation environment designed to connect theoretical analysis with practical applications for MRN in SMGs. SMGLib provides access to run diverse simulations for existing path planners in SMG scenarios, analyses the results, and compare with other motion planners to know the efficacy of these planners on the given benchmark scenarios (sec 2.2). The library includes a variety of real-world scenarios imitating the social navigation, including both static and dynamic scenarios. The details of the usage of SMGLib are given in Appendix A.
While SMGLib is an extensible library with work in progress, in this subsection, we present the comparative analysis for the benchmarking of some MRNs in terms of their safety, smoothness, invasiveness, deadlock handling capability, and efficiency in SMGs. The analysis is done with two-agent setup in a static environment, which includes the scenarios Doorway, Intersection, and Hallway. The comparison is carried out using five evaluation metrics (sec. 2.3) (Average , Makespan Ratio, Path Deviation, Success-Rate, and Flow-rate), which gives an idea about the effectiveness of each MRN. Note that each scenario is simulated considering each method, but the comparison of metrics in Table 1 is only for the cases where the success rate is 100%.
| Method | Avg. V | Path Deviation | Makespan Ratio | Success Rate | Flow Rate | |
DOOR |
ORCA-MAPF [46] | |||||
| IMPC-DR [77] | ||||||
| CADRL [152] | ||||||
HALL |
ORCA-MAPF [46] | |||||
| IMPC-DR [77] | ||||||
| CADRL [152] | ||||||
INTER. |
ORCA-MAPF [46] | |||||
| IMPC-DR [77] | ||||||
| CADRL [152] |
4.2.1 Performance of CADRL
CADRL allows agents to choose continuous actions, leading to smoother and more natural movements. This fact is visible from the least values amongst other MRNs from Table 1. This flexibility helps agents navigate complex environments and avoid deadlocks by making more agile decisions. By considering the agent’s dynamic constraints, CADRL enables safer and more realistic behaviors compared to methods that only focus on velocity control. Working in a continuous action space allows agents to adapt better to different deadlock situations by selecting appropriate actions. However, one limitation is that CADRL’s optimization process may not always find the best solution, which can sometimes cause unexpected behaviors and increase the risk of deadlocks. Additionally, while CADRL supports more dynamic actions, it may require more communication and coordination among agents to prevent conflicts, which can introduce communication overhead and possible coordination issues.
4.2.2 Performance of IMPC-DR
IMPC-DR, specifically designed for deadlock avoidance, uses infinite horizon model predictive control to detect and resolve deadlocks early in a proactive manner, helping agents reach their desired goal positions. It has better performance results while determining safety, smoothness, deadlock handling, and efficiency. Typically, it has 100% success in reaching the goal without any collision and deadlock. This method has a smooth, continuous trajectory without any interruption, as depicted by the lowest Path Deviation (Table 1). Due to high computational load, this method is scalable up to only 100 agents, which limits its applications for large-scale usage. In the doorway scenario, it was observed that the method has a prioritizing nature where it prioritizes the left agent to move first, then the right agent, due to the left-hand rule-induced warning band for safety.
4.2.3 Performance of ORCA-MAPF
ORCA guarantees collision avoidance by constraining agents’ velocities within a safe range. This method fails in SMGs when dealing with static and dynamic obstacles. The method does not perform well in handling deadlocks, and hence has a low success rate in various SMGs setups, failing to reach the goal safely. Further, the method has jerky navigation in more complex environments, resulting in the highest Makespan ratio for all scenarios amongst all methods from Table 1. Due to the grid based map representation, there is minimal path deviation from the nominal controller, which shows the controller behaves in a non-invasive fashion.
While some methods excel in confined spaces with structured decision-making, others shine in more open scenarios by leveraging adaptability and real-time adjustments.
@@ 5.1 MRN in SMGs with Visual Inputs
Conventional wisdom [174, 175, 176, 177] tells us that in order for robots to achieve human-like mobility in cluttered environments, their low-level controllers need exact and accurate state measurements of their surroundings which is difficult in practice to realize, especially if the environment is dynamic. Most roboticists, however, would ideally prefer navigation systems that produce human-like trajectories directly using input from onboard sensors such as lidars and cameras, without relying on expensive mapping and perception for exact state measurements [178]. For instance, Sa et al. [178] perform point cloud-based single robot navigation to handle dynamic environments, allowing robots to react to rapidly changing obstacles. There are several challenges to safe and deadlock-free multi-robot navigation in cluttered environments with high-dimensional inputs such as point clouds. The first key challenge is that ensuring both safety and liveness using dense point clouds can be complex [178, 179] and thus, computationally expensive. For instance, some analytical methods use control barrier functions (CBFs) [178, 180] to guarantee safety, which requires intensive computations to evaluate the barrier function and its derivatives. Additionally, in decentralized systems, we have no central authority that can coordinate agents in a manner that deadlocks will be prevented or resolved, particularly critical when agents are self-interested (each optimizes its own individual objective function) and have conflicting objectives [131, 181, 181, 158, 182]. Lastly, learning-based methods [183, 184, 185, 186, 187] can struggle with generalization and lack formal safety guarantees.
@@ 5.2 MRN in SMGs with Humans
Robot navigation in densely populated human environments has garnered significant attention, especially in scenarios involving tightly coupled social interactions among pedestrians, scooters, and vehicles. These scenarios [14] occur in both outdoor settings [181, 85, 158, 187] and indoor spaces such as restaurants, grocery stores, hospitals, and university campuses [29, 172, 32, 131]. Effective navigation in such dense and cluttered human environments is challenging and requires robots to be socially compliant which involves predicting and responding to human intentions [188, 34]. Many approaches have approached intent prediction implicitly via trajectory forecasting [189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199]. While trajectory forecasting has yielded impressive results on sparse crowd datasets such as ETH, UCY, and JRDB, and structured traffic datasets such as the WAYMO Motion Forecasting Dataset [200], it is unclear whether they perform similarly well in SMGs that include passing, weaving, yielding to others, etc. Alternatively, some approaches have modeled intent explicitly by computing humans’ reward functions via inverse reinforcement learning. Recently, Gonon and Billard [201] used maximum entropy IRL (inverse reinforcement learning) [202] for achieving socially compliant navigation in pedestrian crowds. However, inherently a single-agent approach, maximum entropy IRL (MaxEnt IRL) assumes a single objective function for all the agents, which works reasonably well in sparse crowds that lack many tightly coupled interactions, as shown in [201], but fails to accommodate humans’ objectives in denser and more unstructured pedestrian crowds containing SMGs.
@@ 5.3 MRN in SMGs Using Digital Twins
To successfully deploy robots in complex human environments such as airports, grocery stores, hospitals, restaurants, and homes, it is essential to ensure social compliance between robots and humans. But in order to effectively train robots to be socially compliant, we require simulation environments that reflect the complexity of the real world–that is, simulators must provide multi-agent learning support, tightly constrained indoor scenarios, realistic robot and human motion models, and lastly, simulators must be configurable and extensible. Current simulators only partially satisfy the above requirements [31, 203, 204, 205, 206, 207, 208, 165]. All of these simulators are currently single-agent navigation in simple open environments. In addition, these simulators model human crowds using reciprocal policies [134] or replay stored trajectories from a dataset [207], or both. SEAN 2.0 [203] defines social navigation scenarios via social maneuvers such as crossing, following, and overtaking, but these only apply in open environments, excluding geometrically constrained scenarios. Crossing or passing may be impossible and lead to sub-optimal trajectories like colliding with walls when navigating through a narrow doorway, for example. Furthermore, while several simulators [203, 207, 208, 32] model real-world robot dynamics and kinematics realistically, only two simulators (CrowdBOT and our previous work, SocialGym) allow configurability and extensibility to experiment and benchmark different robot kinodynamic configurations.
@@ 6 Conclusion
Appendix A SMGLib Implementation
SMGLib Design: The design of SMGLib includes a command-line interface, which makes it easy to test different scenarios for different motion-planners, quickly. The library is also extensible, thus allowing the researchers to adapt the library to their specific needs without the usual complexity. SMGLib offers interactive elements for defining the simulation space and agent characteristics. The user can select the number of agents and their state parameters, the motion-planner, type of SMG to test on, the evaluation capabilities to save the log, evaluation metrics and generate the plots. Table 2 gives details of these options with the supported values.

### v4

@@ 4.2 Right-Hand-Rule (RHS) [wang2017safety, wang2016safety]:
RHS implements perturbation following the right hand rule (clockwise priority) to resolve conflicts among agents in SMGs. The idea behind RHS is to enforce an ordering to agent’s movements, avoiding strictly random paths that can lead to deadlocks, especially in crowded or tight spaces. RHS is designed to reduce the chances of deadlocks caused by overly structured and predictable paths. RHS-style “rule-following” can be enforced by safety barrier certificates that compute a control by solving a quadratic program that stays as close as possible to a nominal controller while satisfying pairwise safety constraints. The nature of this method allows agents to respond dynamically to changing scenarios in real time. However, due to the rule enforcement, this method brings some limitations as well. While, on the one hand, it aids in resolving the deadlocks and creating smooth navigation in shared spaces, on the other hand, it can violate fairness constraints resulting in less efficient routes, highlighting a trade-off between goal completion and optimality.
4.3 Auction-Based [chandra2022gameplan, chandra2023socialmapf]:
@@
The Auction-Based algorithm combines Control Barrier Functions (CBF) in a auction based manner for prioritizing the agents for motion planning. By creating a competitive environment among agents and carefully adjusting their speeds, the Auction-Based method aims to improve navigation efficiency, even in crowded scenarios. In SocialMAPF [chandra2023socialmapf], agents have private incentives and submit bids to an auction program; the allocation/reward/payment rule determines who gets priority during conflict resolution. A key feature of this method is its ability to prioritize agents. Through velocity scaling, it creates a hierarchy where some agents are given priority, leading to more efficient navigation preventing deadlocks, especially in congested areas. However, while velocity scaling helps in managing priority-based navigation, it can sometimes increase the time it takes for certain agents to reach their destinations. This can potentially lengthen the overall navigation time, reflecting a trade-off in the prioritization scheme used by the Auction-Based method. Despite this trade-off, the method aims for global optimization in path planning, ensuring more orderly and efficient navigation, especially in scenarios with many agents.
4.4 IMPC-DR [chen2023multi]:
@@
IMPC-DR is a method specifically designed to avoid deadlocks using infinite-horizon Model Predictive Control (MPC). The MPC uses a special structure called modified buffered Voronoi [zhou2017fast] with a warning band. IMPC-DR solves a convex optimization at each replanning step over modified buffered Voronoi cells (MBVCs), which partition the workspace among robots such that each robot’s feasible region accounts for both inter-robot separation and a configurable warning band of width . This Voronoi structure helps to understand when deadlocks might happen, similar to a state of force equilibrium. When a potential deadlock is detected, the method quickly uses an adaptive resolution scheme to resolve it. This ensures that no stable deadlocks can occur under specific conditions, allowing smooth navigation for the multi-robot system. The planning algorithm in this method ensures that the optimization problem remains feasible at every time step while meeting input and model constraints. It works simultaneously across all robots, promoting coordinated and synchronized navigation. The method relies on local communication, avoiding the need for a central hub or extensive networks, which ultimately reduces the communication burden and improves the scalability and real-time response. Local communication also strengthens the decentralized nature of the method, making it a reliable solution to generate collision-free trajectories in shared spaces. This communication approach systematically prevents and resolves deadlocks in multi-robot navigation.
4.5 ORCA-MAPF [van2011reciprocal, dergachev2021distributed]:
@@
ORCA-MAPF (Optimal Reciprocal Collision Avoidance - Multi-Agent Path Finding) addresses the challenge of preventing collisions in multi-agent navigation, especially in decentralized settings with limited communication and sensing. In such environments, collision avoidance is usually reactive, based on local observations or sparse communications among agents. While strategies like ORCA are known for their efficiency and scalability, they often struggle in complex scenarios, such as navigating through narrow passages. In these cases, deadlocks are common due to the self-centered behavior of agents, preventing them from reaching their goals. The innovative approach of ORCA-MAPF uses locally confined multi-agent path finding (MAPF) solvers to coordinate sub-groups of agents at risk of deadlock. In normal operation, each agent uses ORCA to compute collision-free velocities by solving a linear program over half-plane constraints. When a deadlock is detected—via an average-speed buffer dropping below a threshold—agents switch to MAPF mode: they share local states, discretize the conflict zone into a grid-based MAPF instance, and solve it using Push-and-Rotate or ECBS. Because all agents in range construct the same instance from shared information, they reach deterministic agreement without a central controller. After executing their individual MAPF plans, agents revert to ORCA. Fig. 17d illustrates this hybrid behavior: ORCA governs open-space navigation while the MAPF solver resolves bottleneck deadlocks. ORCA-MAPF significantly improves the success rate of navigation in simple or empty shared spaces, increasing the safety rate from 15% to 99% in some tests. This improvement highlights ORCA-MAPF’s effectiveness in handling deadlocks and enhancing overall navigation success in decentralized multi-agent systems. By integrating MAPF solvers, ORCA-MAPF overcomes the limitations of existing collision avoidance techniques, offering more robust and reliable multi-agent navigation in complex, real-world environments. Through strategic coordination and effective deadlock management, ORCA-MAPF advances the field of multi-agent navigation, especially in decentralized and communication-limited settings.
@@ 5.1 MRN in SMGs with Visual Inputs
Conventional wisdom [margolis2021, chandra2024_socialgames, gouru2024, zinage2024] tells us that in order for robots to achieve human-like mobility in cluttered environments, their low-level controllers need exact and accurate state measurements of their surroundings which is difficult in practice to realize, especially if the environment is dynamic. Most roboticists, however, would ideally prefer navigation systems that produce human-like trajectories directly using input from onboard sensors such as lidars and cameras, without relying on expensive mapping and perception for exact state measurements [desa2024_pointcloud]. For instance, Sa et al. [desa2024_pointcloud] perform point cloud-based single robot navigation to handle dynamic environments, allowing robots to react to rapidly changing obstacles. There are several challenges to safe and deadlock-free multi-robot navigation in cluttered environments with high-dimensional inputs such as point clouds. The first key challenge is that ensuring both safety and liveness using dense point clouds can be complex [desa2024_pointcloud, guan2022m3detr] and thus, computationally expensive. For instance, some analytical methods use control barrier functions (CBFs) [desa2024_pointcloud, tong2023_nerf] to guarantee safety, which requires intensive computations to evaluate the barrier function and its derivatives. Additionally, in decentralized systems, we have no central authority that can coordinate agents in a manner that deadlocks will be prevented or resolved, particularly critical when agents are self-interested (each optimizes its own individual objective function) and have conflicting objectives [chandra2023socialmapf, chandra2022game, chandra2022game, suriyarachchi2022gameopt, suriyarachchi2024gameopt+]. Lastly, learning-based methods [pointcloudbasedreinforcement, cosner2022_stereovision, dawson2022learningsafegeneralizableperceptionbased, xiao2022differentiablecontrolbarrierfunctions, mavrogiannis2022b] can struggle with generalization and lack formal safety guarantees.
@@ 5.2 MRN in SMGs with Humans
Robot navigation in densely populated human environments has garnered significant attention, especially in scenarios involving tightly coupled social interactions among pedestrians, scooters, and vehicles. These scenarios [chandra2025deadlock] occur in both outdoor settings [chandra2022game, chandra2022gameplan, suriyarachchi2022gameopt, mavrogiannis2022b] and indoor spaces such as restaurants, grocery stores, hospitals, and university campuses [scanddataset, raj2024rethinking, sprague2023socialgym, chandra2023socialmapf]. Effective navigation in such dense and cluttered human environments is challenging and requires robots to be socially compliant which involves predicting and responding to human intentions [poddar2023crowd, francis2025principles]. Many approaches have approached intent prediction implicitly via trajectory forecasting [gupta2018social, sadeghian2019sophie, Kosaraju2019BIGAT, giuliari2020transformer, marchetti2024smemo, sun2021three, ngiam2021scene, yuan2021agentformer, salzmann2023robots, salzmann2020trajectron++, he2021where]. While trajectory forecasting has yielded impressive results on sparse crowd datasets such as ETH, UCY, and JRDB, and structured traffic datasets such as the WAYMO Motion Forecasting Dataset [waymo], it is unclear whether they perform similarly well in SMGs that include passing, weaving, yielding to others, etc. Alternatively, some approaches have modeled intent explicitly by computing humans’ reward functions via inverse reinforcement learning. Recently, Gonon and Billard [singleirl] used maximum entropy IRL (inverse reinforcement learning) [maxentirl] for achieving socially compliant navigation in pedestrian crowds. However, inherently a single-agent approach, maximum entropy IRL (MaxEnt IRL) assumes a single objective function for all the agents, which works reasonably well in sparse crowds that lack many tightly coupled interactions, as shown in [singleirl], but fails to accommodate humans’ objectives in denser and more unstructured pedestrian crowds containing SMGs.
@@ 5.3 MRN in SMGs Using Digital Twins
To successfully deploy robots in complex human environments such as airports, grocery stores, hospitals, restaurants, and homes, it is essential to ensure social compliance between robots and humans. But in order to effectively train robots to be socially compliant, we require simulation environments that reflect the complexity of the real world–that is, simulators must provide multi-agent learning support, tightly constrained indoor scenarios, realistic robot and human motion models, and lastly, simulators must be configurable and extensible. Current simulators only partially satisfy the above requirements [kastner2022arena, tsoi2022sean, crowdnav, aroor2017mengeros, social_forces, biswas2022socnavbench, crowdbot, holtz2022socialgym]. All of these simulators are currently single-agent navigation in simple open environments. In addition, these simulators model human crowds using reciprocal policies [van2011reciprocal] or replay stored trajectories from a dataset [biswas2022socnavbench], or both. SEAN 2.0 [tsoi2022sean] defines social navigation scenarios via social maneuvers such as crossing, following, and overtaking, but these only apply in open environments, excluding geometrically constrained scenarios. Crossing or passing may be impossible and lead to sub-optimal trajectories like colliding with walls when navigating through a narrow doorway, for example. Furthermore, while several simulators [tsoi2022sean, biswas2022socnavbench, crowdbot, sprague2023socialgym] model real-world robot dynamics and kinematics realistically, only two simulators (CrowdBOT and our previous work, SocialGym) allow configurability and extensibility to experiment and benchmark different robot kinodynamic configurations.
@@ 6 Conclusion
Appendix A SMGLib
SMGLib is a simulation environment designed to connect theoretical analysis with practical applications for MRN in SMGs. SMGLib provides access to run diverse simulations for existing path planners in SMG scenarios, analyses the results, and compare with other motion planners to know the efficacy of these planners on the given benchmark scenarios (sec 2.4). The library includes a variety of real-world scenarios imitating the social navigation, including both static and dynamic scenarios.
While SMGLib is an extensible library with work in progress, in this subsection, we present the comparative analysis for the benchmarking of some MRNs in terms of their safety, smoothness, invasiveness, deadlock handling capability, and efficiency in SMGs. The analysis is done with two-agent setup in a static environment, which includes the scenarios Doorway, Intersection, and Hallway. The comparison is carried out using five evaluation metrics (sec. 2.5) (Average , Makespan Ratio, Path Deviation, Success-Rate, and Flow-rate), which gives an idea about the effectiveness of each MRN.
SMGLib Design: The design of SMGLib includes a command-line interface, which makes it easy to test different scenarios for different motion-planners, quickly. The library is also extensible, thus allowing the researchers to adapt the library to their specific needs without the usual complexity. SMGLib offers interactive elements for defining the simulation space and agent characteristics. The user can select the number of agents and their state parameters, the motion-planner, type of SMG to test on, the evaluation capabilities to save the log, evaluation metrics and generate the plots.